### 02 Apr 2026 v2

# Random Stability of Random Variables

Abstract.
For a random variable $N=0,1,2,\ldots$ we study the following question: When does the sum of $N$ many independent and identically distributed copies of a random variable $X$ have the same law a a nontrivial rescaling of $X$? We show that such $N$-stable random variable exists if and only $1<\mathbb{E}[N]<\infty$. Under an additional assumption $\mathbb{E}[N\ln N]<\infty$, we describe all $N$-stable $X$. We also study a converse problem: For a given $X\geq 0$ with $\mathbb{E}[X]=1$, we study the set of all $N$ such that $X$ is $N$-stable. Distributions of $N$ form a semigroup with respect to composition of probability generating functions. We show these probability generating functions need to commute with respect to composition. We present explicit families of composition semigroups. Equivalent formulations have appeared in difference forms, and this article aims to unify and extend them.
2020 Mathematics Subject Classification:
60E07, 60E10, 60J80. Branching processes, stable distributions, strong stability, characteristic function, Linnik distribution, Mittag-Leffler distribution, Yule process, Poincare functional equation.
1. Introduction
1.1. The main definition
The topic of this article is the concept called random stability or $N$-stability.
Definition 1.
Take a random variable $N$ with values $0,1,2,\ldots$ with distribution $P$. A probability measure $Q$ on $\mathbb{R}$, or, equivalently, a random variable $X\sim Q$, is called $N$-stable or, equivalently, $P$-stable, if there exists a $c>0$ such that the following equality in law holds:
(1) $X_{1}+\ldots+X_{N}\stackrel{{\scriptstyle d}}{{=}}cX,$
where $X_{1},X_{2},\ldots\sim Q$ are independent copies of $X$, independent of $N$. of independent identically distributed random variables, independent of $N$.
Alternatively, the article [38] uses the term branching stability to stress connection with branching processes. We will see in our article that this connection is indeed critical. In [25, Definition 3], the term strictly $N$-stable is used.
Definition 1 is a generalization of a well-known concept of strictly stable distributions $Q$, which are $N$-stable in terms of this Definition 1 for a constant $N=n$ for each $n=2,3,\ldots$ The theory of strictly stable distributions is classic, see [45, 51].
Importantly, we can rewrite the main equation from Definition 1. Taking the characteristic function (Fourier transform) $f(u)=\mathbb{E}[e^{iuX}]$ of $X$, and the probability generating function $\varphi$ of $N$: $\varphi(s)=\mathbb{E}[s^{N}]$, we rewrite this as the Poincare functional equation:
(2) $\varphi(f(u))=f(cu),\quad u\in\mathbb{R}.$
If $X\geq 0$, then the same equation holds for the Laplace instead of the Fourier transform.
1.2. Motivation
For random stability, there has been a lot of research starting from the 1990s, as well as some earlier research in relation to the classic theory of branching processes. See [29] for a survey of geometric stable distributions ($N$-stable for geometric $N$), and a more detailed description in [23, Chapter 9]. A classic application for classic stable and random stable laws is asset returns, see [29, Remark 2.2] and references therein; also [22, Chapter 4], page 157. For $N$-stable random variables with general $N$, see [22, Chapter 4].
However, existing literature is fragmented. Equivalent formulations have appeared in difference forms, and this article aims to unify and extend them. It contains exposition, new research, and critique of existing research. Theorems 1,  2, and 3 are our main results. We could not find Theorem 1 in other articles. Theorem 2 was shown in [30] but under special assumptions. Theorem 3 was shown in special cases, see [13].
1.3. Organization of the article
Section 2 contains main results: Theorems 12, and 3, as well as most lemmas. In Section 3, we fix a distribution $Q$ on the real line and study $G(Q)$, which is the set of distributions $P$ on $0,1,2,\ldots$ for which $Q$ is $P$-stable. In Section 4, we discuss two different but equivalent definitions of random stability. Section 5 discusses several cases of such semigroups. Section 6 discusses the case of commuting semigroups of probability generating functions equivalent to continuous-time branching processes. The last two sections are devoted to proofs and conclusions.
2. Main Results
2.1. Existence results
Theorem 1 is the main result of this article. As mentioned earlier, we could not find it in the literature, despite its importance. The closest was [35, Corollary 1.2] with (in their notation) $A_{i}=1/c$. This discusses a generalization of random stability to so-called it smoothing problems. But their proof applies only to nonnegative $N$-stable $X$.
Theorem 1.
Assume $N$ is not identically zero or one. There exists an $N$-stable random variable $X$ which is not identically zero if and only if $1<\mathbb{E}[N]<\infty$.
The following is a well-known uniqueness result, which can be found in the classic monograph on branching processes: [1, Chapter I], Section 10, Theorem 2.
Proposition 1.
Assume $\mathbb{E}[N]>1$ and $\mathbb{E}[N\ln N]<\infty$. Then an $N$-stable $X\geq 0$ is unique in law under condition $\mathbb{E}[X]=1$. Then $c=\mathbb{E}[N]$ in (1).
An earlier monograph [17, Chapter 1], Theorem 7.3, has this result for the case $\mathbb{E}[N^{2}]<\infty$. The following result was proved in [1, Chapter I], Section 10, Theorem 3. Let $(N_{k})$ be a branching process with $N$ offsprings and $c=\mathbb{E}[N]\in(1,\infty)$. This case is called supercritical. There exists a sequence $(C_{k})$ of positive numbers such that
(3) $\frac{C_{k+1}}{C_{k}}\to c,\quad\frac{N_{k}}{C_{k}}\stackrel{{\scriptstyle a.s.}}{{\to}}X,$
where $X\geq 0$ is an $N$-stable random variable.
Remark 1.
There are two cases, when $c=\mathbb{E}[N]\in(1,\infty)$; see also discussion in [5]. For a conceptual proof, see [37]. For other proofs, see [47, Theorem 3.1], [48, Theorem 4.3].
  • If $\mathbb{E}[N\ln N]<\infty$, we can simply take $C_{k}=c^{k}$. Then $\mathbb{E}[X]=1$.
  • If $\mathbb{E}[N\ln N]=\infty$, we need to take $C_{k}\neq c^{k}$, and $\mathbb{E}[X]=\infty$.
Definition 2.
We call such $X$ from Remark 1 the standard $N$-stable random variable, and the distribution of $X$ is called the standard $N$-stable distribution.
2.2. Strictly stable distributions on the real line
Recall a well-known classification of stictly stable distributions $Q$, see [45, Chapter 1] or [51, Chapter V]. Its characteristic function (Fourier transform)
$\int_{\mathbb{R}}e^{\mathrm{i}ux}\,Q(\mathrm{d}x)=e^{-g(u)},\quad\mathrm{i}=\sqrt{-1},$
can be represented as
(4) $g(u)=(\beta+\gamma\mathrm{i}\mathrm{sgn}(u))|u|^{\alpha}$
with $\alpha\in(0,2]$, $\beta>0$, and the following additional restrictions:
  • if $\alpha=2$ then $\gamma=0$ and $Q=\mathcal{N}(0,2\beta)$;
  • if $\alpha=1$ then $\gamma$ is unrestricted and $Q$ is shifted Cauchy;
  • if $\alpha\in(0,1)\cup(1,2)$ then $|\gamma|\leq|\mathrm{tg}(\pi\alpha/2)|$.
Of course, we also have the trivial case $\beta=\gamma=0$, when $Q=\delta_{0}$. Denote the set of such functions $g$ as $\mathcal{G}$. Such distribution $Q$ is also called strictly $\alpha$-stable to stress its dependence upon the index $\alpha$. Note that the function $g$ satisfies
(5) $g(cu)=c^{\alpha}g(u),\quad c>0,\quad u\in\mathbb{R}.$
A standard reference is the classic monograph [45, Chapter 1]: Equivalent Definitions 1.1.1, 1.1.4, 1.1.5, 1.1.6 of stable distributions, and Properties 1.2.6, 1.2.8 of strictly stable distributions. Also, see other monographs: [36, Chapter 3]: Theorem 5.7.3 and discussions after this; a more recent book [51, Chapter V]: Section 7 for general results about strict stability; Theorem 7.6 for strictly stable symmetric distributions; Theorem 3.5 for positive strictly stable symmetric distributions; Theorem 3.5 for positive strictly stable random variables. Finally, we point out to [13, Chapter 4], section 4.5, page 130. We stress that much of classic research involves stable not strictly stable distributions, which arise as weak limits
(6) $\frac{X_{1}+\ldots+X_{n}}{a_{n}}-b_{n},\quad n\to\infty,$
for independent identically distributed $X_{1},X_{2},\ldots$ But we use strictly stable distributions here in this article, with $b_{n}=0$.
2.3. Classification of $N$-stable $X$ for given NN
In Theorem 2, we classify $N$-stable distributions. This replicates the results [30, Proposition 2.1, Proposition 3.1] but without an artificial additional assumption in Remark 2 below. The $N$-stable laws are represented via a product involving a standard $N$-stable random variable, and an independent strictly $\alpha$-stable random variable.
Theorem 2.
Assume $\mathbb{E}[N]>1$ and $\mathbb{E}[N\ln N]<\infty$. Any $N$-stable $X$ can be represented as the product
(7) $X\stackrel{{\scriptstyle d}}{{=}}Y^{1/\alpha}Z,$
where $Y$ is the standard $N$-stable random variable, and $Z$ is a strictly $\alpha$-stable random variable independent of $Y$, for some $\alpha\in(0,2]$. The characteristic function of $X$ can be represented as
(8) $f(u)=\mathbb{E}\left[e^{\mathrm{i}uX}\right]=L(g(u)),$
where $L(u)=\mathbb{E}[e^{-uY}]$ is the Laplace transform of this standard $N$-stable $Y$, and $g\in\mathcal{G}$ is such that the characteristic function of $Z$ is $e^{-g}$. Moreover, if $X$ is $N$-stable with index $\alpha$, then it satisfies (1) with
(9) $c=(\mathbb{E}[N])^{1/\alpha}.$
Separately, we state an important uniqueness lemma.
Lemma 1.
Assume $L_{1}$ and $L_{2}$ are Laplace transforms of two probability measures $Q_{1}$ and $Q_{2}$ on the positive half-line with mean $1$. Let $g_{1},g_{2}\in\mathcal{G}$ be such that $L_{1}\circ g_{1}\equiv L_{2}\circ g_{2}$. Then $L_{1}\equiv L_{2}$ and $g_{1}\equiv g_{2}$.
2.4. Removal of artificial assumptions
This result is important for the following subtle reason. Previously, such representation result was proved by [30] under an additional condition on the Laplace transform $L(u)=\mathbb{E}[e^{-uY}]$ of this standard $N$-stable $Y$.
Remark 2.
For any two infinitely divisible characteristic functions $e^{-f}$ and $e^{-g}$ on the real line, $L\circ f\equiv L\circ g$ implies $f\equiv g$.
This condition was first introduced in [53] and used for the converse of the transfer theorem. This theorem is about a transfer from weak limits of sums of independent random variables to weak limits of random sums of these variables, [13, Chapter 4]: Section 4.1, Theorem 4.1.2, and references therein. Says [53]: The following condition is presumably true, but so far no one was able to prove or disprove it. It turns out from our research that we do not, in fact, need to prove this condition! Indeed, from 1973 until now, to the best of our knowledge, no one could prove this in the general form. There were many discussions of this condition in a recent book [13], which is an unabridged republication from 1996: [13, Chapter 4], Section 2. To quote: No one has managed either to prove this or to refute it yet. There, it was proved in several cases. In [23, Chapter 8], it was mentioned but unproved in Remark 8.8.16.
2.5. Examples of random stable variables
Here, we consider the three cases of $N$-stable distributions: (a) symmetric; (b) Gaussian; (c) on the half-line.
Definition 3.
The random variable $X$ (or its distribution $Q$) is called a symmetric $N$-stable if it is $N$-stable and $X\stackrel{{\scriptstyle d}}{{=}}-X$.
Lemma 2.
Assume $1<\mathbb{E}[N]$ and $\mathbb{E}[N\ln N]<\infty$. Any symmetric $N$-stable $X$ has characteristic function $f(u)=L(\beta|u|^{\alpha})$ for some $\beta>0$ and $\alpha\in(0,2]$, where $L$ is the Laplace transform of the standard $N$-stable random variable $Y$. This $X$ can be represented as
$X\stackrel{{\scriptstyle d}}{{=}}Y^{1/\alpha}Z,\quad\mathbb{E}\left[e^{\mathrm{i}uZ}\right]=\exp\left(-\beta|u|^{\alpha}\right).$
Here, $Z$ is independent of $Y$ and is symmetric $\alpha$-stable.
Definition 4.
Any $N$-stable random variable with index $\alpha=2$ is called a Gaussian $N$-stable random variable.
We stress its similarity with classic Gaussian random variables. In the literature, sometimes a Gaussian $N$-stable $X$ is defined using (1) with (9) with $\alpha=2$. An additional restriction $\mathbb{E}[X^{2}]=\infty$ is introduced in [24, Definition 2]. Confusingly, the opposite restriction $\mathbb{E}[X^{2}]<\infty$ was introduced in [13, Chapter 4], see Definition 4.6.1; or [25, Definition 2], or [23, Chapter 8], Definition 8.3.2. But we do not need these restrictions in the definition of a Gaussian $N$-stable random variable, as seen in the following lemma.
Lemma 3.
Assume $\mathbb{E}[N]>1$ and $\mathbb{E}[N\ln N]<\infty$. Any Gaussian $N$-stable $X$ is symmetric: $X\stackrel{{\scriptstyle d}}{{=}}-X$. It satisfies $\mathbb{E}[X^{2}]<\infty$. Taking $Y$ the standard $N$-stable random variable, with Laplace transform $L$, we can represent $X=\sqrt{Y}Z$, where $Z\sim\mathcal{N}(0,\sigma^{2})$ is independent of $Y$. The characteristic function of $X$ is given by $\mathbb{E}\left[e^{\mathrm{i}uX}\right]=L(\sigma^{2}u^{2}/2)$.
The final case is a nonnegative $N$-stable random variable $X$. It is easier to study their Laplace transforms than characteristic functions.
Lemma 4.
If $\mathbb{E}[N]>1$ and $\mathbb{E}[N\ln N]<\infty$, any nonnegative $N$-stable $X$ has Laplace transform $\mathbb{E}\left[e^{-uX}\right]=L(\beta u^{\alpha})$ for the Laplace transform $L$ of the standard $N$-stable random variable $Y$, and for some $\alpha\in(0,1]$ and $\beta>0$. This $X$ can be represented as $X\stackrel{{\scriptstyle d}}{{=}}Y^{1/\alpha}Z$, where $Z$ is the stable subordinator of index $\alpha$, independent of $Y$, defined via Laplace transform
(10) $\mathbb{E}\left[e^{-uZ}\right]=e^{-\beta u^{\alpha}},\quad u>0.$
Remark 3.
A nonnegative $N$-stable $X$ can have only index $\alpha\leq 1$, not $\alpha\in(1,2]$.
2.6. Geometric stability
The most classic example is for the geometric $N_{p}$ for $p\in(0,1)$, with distribution
(11) $\mathbb{P}(N_{p}=n)=p(1-p)^{n-1},\quad n=1,2,\ldots;$
and with expected value and the probability generating function
(12) $\mathbb{E}[s^{N}]=\frac{ps}{1-(1-p)s},\quad\mathbb{E}[N_{p}]=\frac{1}{p}.$
Then the standard $N_{p}$-stable distribution of every $p\in(0,1)$ is the standard exponential with mean $1$: It has density and the Laplace transform
(13) $p(x)=e^{-x},\quad x\geq 0;\quad L(u)=\mathbb{E}[e^{-uY}]=\frac{1}{1+u},\quad u\geq 0;\quad Y\sim\mathrm{Exp}(1).$
A Gaussian $N_{p}$-stable distribution is also well-known: The Laplace distribution with density and the characteristic function
$p(x)=0.5\beta e^{-\beta|x|},\quad x\in\mathbb{R};\quad f(u)=\frac{1}{1+\beta^{2}u^{2}}.$
More generally, a symmetric $N_{p}$-stable distribution is known: This is the Linnik distribution with characteristic function
$f(u)=\frac{1}{1+\beta|u|^{\alpha}}.$
with $0<\alpha\leq 2$. See also [23, Chapter 9], Subsection 9.4.2.1, Theorem 9.4.6. This distribution is absolutely continuous on the real line, but its density is explicitly known only for $\alpha=2$, when it becomes the Laplace distribution. The positive $N_{p}$-stable distribution has Laplace transform with $\beta>0$ and $\alpha\in(0,1]$:
$L(u)=\frac{1}{1+\beta u^{\alpha}}.$
This is commonly known as the Mittag-Leffler distribution; see [23, Chapter 9], Subsection 9.4.2.2, Theorem 9.4.9. Also, in [13, Chapter 2], these were named Kovalenko distributions, see (3.2) or page 30. This distribution is also absolutely continuous, but its density is explicitly known only for $\alpha=1$ or $\alpha=0.5$. In the former case, this distribution is exponential, as in (13). In the latter case, for $\beta=1$, it has density
$f(x)=\frac{1}{\sqrt{\pi x}}-\frac{2e^{x}}{\sqrt{x}}\int_{\sqrt{x}}^{\infty}e^{-z^{2}}\,\mathrm{d}z.$
These were the earliest and most studied examples. An entire chapter [23, Chapter 9] is devoted to these geometric stable distributions. See also a great survey [29] with large bibliography. Such geometric stable distributions were studied as far back as in 1984 in [26] and also in one early article by T. Kozubowski [28]. Further representation results for the Linnik and the Mittag-Leffler distributions are given in [27].
2.7. The infinitely divisible case
The following example is well known, see [17, Chapter I, Example 8.5; Chapter V, Example 13.2], [38, Section 8], [24, Section 6], Exampples 2 and 4, [7, Example 1]; [42, Example 3], and physical applications in [16, Section 8]. Here $X$ is a Gamma random variable with shape $1/k$ and certain rate parameter $\beta>0$. We remind the readers that it has Laplace transform
(14) $\mathbb{E}[e^{-uX}]=(1+\beta u)^{-1/k},\quad u>0.$
and $N$ has probability generating function for some $p\in(0,1)$:
(15) $\varphi(s)=\frac{p^{1/k}s}{(1-(1-p)s^{k})^{1/k}}.$
And $X$ is $N$-stable for any $m>1$. This is related to the classic geometric-exponential pair from (13). Indeed, since the sum of $k$ i.i.d. Gamma from (14) is exponential:
$X_{1}+\ldots+X_{k}\sim\mathrm{Exp}(1).$
And the probability generating function from (15) corresponds to the random variable $kM$, where $M$ is negative binomial with shape $1/k$, and has probability generating function
$\mathbb{E}[s^{M}]=\left(\frac{ps}{1-(1-p)s}\right)^{1/k}$
The sum of $k$ independent identically distributed copies of this $M$ is, in fact, geometric, it is distributed as $N_{p}$ from (11):
$M_{1}+\ldots+M_{k}\stackrel{{\scriptstyle d}}{{=}}N_{p}.$
In Lemma 5, we generalize this for general infinitely divisible distributions. We remind the readers the classic definition. For background, see [51].
Definition 5.
A distribution $Q$ on the real line, or, equivalently, a random variable $X$, is called infinitely divisible if for each $k=1,2,\ldots$, we can represent $X$ as a sum of $k$ independent identically distributed random variables $X_{1},\ldots,X_{k}$:
$X\stackrel{{\scriptstyle d}}{{=}}X_{1}+\ldots+X_{k}.$
In terms of characteristic functions $f(u)$, or Laplace transforms $L(u)$, or probability generating functions $\varphi(s)$, a necessary and sufficient condition is that for every $k=2,3,\ldots$ the $k$th root of this function: $f^{1/k}(u)$, $L^{1/k}(u)$, $\varphi^{1/k}(u)$ must also be a characteristic function, or a Laplace transform, or a probability generating function of some distribution.
Lemma 5.
Take infinitely divisible random variables $N=0,1,2,\ldots$ and $X$ on the real line. Pick a $k=2,3,\ldots$ and decompose
$N=M_{1}+\ldots+M_{k},\quad X=Y_{1}+\ldots+Y_{k}$
into $k$ independent copies of $M$ and $Y$, respectively. Then $X$ is $N$-stable if and only if $Y$ is $kM$-stable.
2.8. Commuting distributions
The uniqueness result in Lemma 1 for the representation from Theorem 2 allows us to prove the following key statement. First, we consider the composition operation
(16) $\varphi\circ\psi(s)\equiv\varphi(\psi(s))$
for the two probability generating functions $\varphi$ and $\psi$, or, equivalently, the corresponding distributions. This operation is clearly associative, so the set of all probabiltiy distributions on $\{0,1,2,\ldots\}$ is a semigroup under this operation. In terms of the random variables $M$ and $N$ with probability generating functions $\varphi$ and $\psi$, their composition $\varphi\circ\psi$ is also a probability generating function of the random variable $N_{1}+\ldots+N_{M}$, where $N_{1},N_{2},\ldots$ are independent (of each other and $M$) copies of $N$. The unit element for this composition operation is $\varphi(s)=s$, corresponding to the variable $M=1$.
Definition 6.
We say that the probability generating functions $\varphi$ and $\psi$, or their corresponding distributions, or random variables $M$ and $N$, commute if
$\varphi\circ\psi\equiv\psi\circ\varphi.$
The key result is below. It was shown in other literature in particular cases, but not in this general form. For such results for or Gaussian $N$-stable random variables, see [13, Chapter 4], Theorem 4.6.1 and Corollary 4.6.1; [25, Theorem 1], [23, Chapter 8], Theorem 8.3.4 and references therein; [42, Theorem 2.1]. The property of commutative semigroups was also shown in [29, Remark 2.4]. This survey [29] is very comprehensive for 1990s. However, they claim that geometric family is unique among such explicit commuting families. We discuss in Section 2 that this is not true.
Theorem 3.
Assume $\mathbb{E}[M]<\infty$ and $\mathbb{E}[N\ln N]<\infty$. If $X$ is $N$-stable, then $M$ and $N$ commute if and only if $X$ is $M$-stable. In this case, $\mathbb{E}[M\ln M]<\infty$. The standard $N$-stable and the standard $M$-stable distributions are the same. The index $\alpha$ and the representation of $X$ from Theorem 2 for both $M$ and $N$ are the same.
Remark 4.
As a corollary of Theorem 3, we get that $M$ and $N$ commute if and only if their standard random stable distributions coincide.
Lemma 6.
Assume $\mathbb{E}[N\ln N]<\infty$. If $X$ and $Y$ are $N$-stable, and $X$ is $M$-stable, then $Y$ is also $M$-stable. In this case, $\mathbb{E}[M\ln M]<\infty$, and $\mathbb{E}[N]>1$, and $\mathbb{E}[M]>1$.
3. Composition Semigroups
3.1. Introduction to the problem
Fix a probability measure $Q$ on the real line. In this section, we give an overview of known results and add a few new ones on the following topic: Find the set of all distributions $P$ on the nonnegative integers such that $Q$ is $P$-stable. Let $N\sim P$ and assume $\mathbb{E}[N\ln N]<\infty$. Then from Theorem 3 and Lemma 6, it is sufficient to consider $Q$ which is the standard $N$-stable for any such $N$. This allows to use the Laplace transform $L$ of $Q$. Such $X\sim Q$ satisfies $X\geq 0$ and $\mathbb{E}[X]=1$.
However, we stress that not every $X\geq 0$ can have nontrivial $N$ such that $X$ is $N$-stable. First, to be $N$-stable for at least one $N$ with $1<\mathbb{E}[N]<\infty$, the distribution of $X$ must be absolutely continuous on $(0,\infty)$; see [1, Chapter I], Section 10, Theorem 4; Secton 12, Corollary 1; also [17, Chapter 1], Theorem 8.3. However, this absolute continuity is very far from a sufficient condition for existence of $N$ such that $Q$ is $N$-stable. Recall the pair of negative binomial $N$ and Gamma $X$ random variables from (14) and (15) such that $X$ is $N$-stable. Compare it with Lemma 7.
Lemma 7.
A Gamma random variable with shape parameter $\alpha\neq 1/n$ for $n=1,2,\ldots$ is not $N$-stable for any $N$.
The idea of the proof of Lemma 7 is to consider the behavior of the probability generating function of $N$ at $s=0$. From (2), which involves the behavior of the Laplace transform of $X$ at infinity. This approach is useful to disprove that $\varphi$ is a probability generating function in other cases. Such distribution $Q$ might have an atom at 0, that is, have $Q({0})=\mathbb{P}(X=0)>0$; and this happens if and only if $N$ has an atom at zero: $\mathbb{P}(N=0)>0$. By the way, the case of an atom at zero can be reduced to the case of no atom at zero. See [1, Chapter I], Section 12; [38, Section 9]; [16, Section 3], Definitions 3.1, 3.2; Theorem 3.2.
3.2. Analytic description
Let us describe such $N$ analytically. Let $\pi=\mathbb{P}(X=0)$. The Laplace transform $L$ is a one-to-one strictly decreasing function $0,\infty)\to(\pi,1]$, and we can find its inverse $L^{\leftarrow}$ (denoted using the arrow to distinguish it from $1/L$):
(17) $L(L^{\leftarrow}(s))\equiv s,\quad s\in(\pi,1].$
Recall that the functional equation (2) holds for the Laplace transform $L$ of $X$ as well as for its characteristic function $f$:
(18) $\varphi(L(u))=L(cu),\quad u\geq 0.$
Using (17), we rewrite (18) as Schroder functional equation:
$L^{\leftarrow}(\varphi(s))=cL^{\leftarrow}(s),\quad s\in(\pi,1].$
In complex analysis, Schroder’s equation is well known, see [49] for a bibliography and a survey of results. We can also rewrite (18) using (17) as follows:
(19) $\varphi(s)=L(cL^{\leftarrow}(s)),\quad s\in(\pi,1].$
Equation (19) is the key formula in a seminal article [7] on random stability. It will play an important role in this article as well.
This defines the probability generating function on the interval $(\pi,1]\subseteq[0,1]$. Maybe this is not the entire $[0,1]$, but even if this interval is smaller, it is large enough to uniquely determine the function $\varphi$. Indeed, this function $\varphi$, as any probability generating function, is analytic in the unit disc $\mathbb{D}:=\{z\in\mathbb{C}\mid|z|<1\}$. Therefore, the values of $\varphi$ on this interval uniquely determine it: This is a classic result in complex analysis.
3.3. Semigroup definitions and properties
Take any probability measure $Q$ on the half-line with mean $1$. We can define two semigroups:
  • The set $G(Q)$, or, equivalently, $G(X)$, of all probability generating functions $\varphi$ (or, equivalently, all distributions $P$ on $\{0,1,2,\ldots\}$) such that $Q$ is $P$-stable.
  • The set $S(X)$ or $S(Q)$ of all $c\geq 1$ such that the right-hand side of (19) is a probability generating function.
For a given $Q$, one of the two possibilities exist. Either there is no $N$ except $N\equiv 1$ such that (1) holds. In this case, $G(Q)$ contains only the identity probability generating function $\varphi(s)\equiv s$; and $S(Q)=\{1\}$. Or $Q$ is $N$-stable for some nontrivial $N$. The main formula (19) defines a mapping $\mathcal{M}:S(Q)\to G(Q)$ as follows:
(20) $\mathcal{M}:c\mapsto L(cL^{\leftarrow}(\cdot))$
Lemma 8.
The set $G(Q)$ is closed under composition and weak convergence. The set $S(Q)$ is closed under multiplication and under the usual convergence of real numbers. The mapping $\mathcal{M}$ from (20) is a bijection, and it preserves this semigroup operation and convergence. If $1$ is the limit point of $S(Q)$, then $S(Q)=[1,\infty)$.
Below we state an important result about finite moments for distributions in $G(Q)$.
Lemma 9.
Pick a distribution $Q$ on the real line, and assume $G(Q)\neq\{1\}$ is nontrivial. Then let $N$ be any nontrivial random variable with distribution from $G(Q)$, and let $Y$ be the standard $N$-stable random variable. Then:
  • $\mathbb{E}[N\ln N]<\infty$ if and only if $\mathbb{E}[Y]<\infty$;
  • $\mathbb{E}[N\ln^{a+1}N]<\infty$ for a fixed $a>0$ if and only if $\mathbb{E}[Y\ln^{a}Y]<\infty$;
  • $\mathbb{E}[N^{k}]<\infty$ for a fixed $k=2,3,\ldots$ if and only if $\mathbb{E}[Y^{k}]<\infty$.
As a corollary, if one distribution $P\in G(Q)$ satisfies $\mathbb{E}[N\ln^{a+1}N]<\infty$ for a fixed $a\geq 0$, or $\mathbb{E}[N^{k}]<\infty$ for a fixed $k=2,3,\ldots$ where $N\sim P$, then all other distributions in $G(Q)$ satisfy this property, too.
3.4. Parameterization
Here, we use $c=\mathbb{E}[N]\geq 1$ (for the standard $X$, we have $\alpha=1$), and the semigroup operation on $c$ is multiplication. This exactly matches the notation in [7] with $e^{t}$ instead of $c$, and $t\geq 0$. The corresponding semigroup operation here is addition. However, in other literature, a different parameterization is used: $p\in(0,1)$. Historically, $N$-stable $X$ arose as weak limits of scaled random sums of random variables. Much like classic stable distributions are weak lmits of scaled and shifted sums of a deterministic nut growing number of independent identically distributed random variables. Therefore, the notation was used is $\nu_{p}$ instead of $N$, with
(21) $p\nu_{p}\stackrel{{\scriptstyle d}}{{\to}}Z\geq 0,\quad p\in\Xi\subseteq(0,1),\quad p\downarrow 0;$
where $Z$ is, in fact, the standard $\nu_{p}$-stable random variable for any $p$; and $\Xi$ is a subset of $(0,1)$ with limit point 0. They found it more convenient to impose this limit assumption, and not derive it from scratch. Often, another additional condition was imposed: $\mathbb{E}[\nu_{p}]=1/p$. Normalizing by convergence and expectation at the same time. But this might not be conistent, and we avoid this notation here. The reason is subtle: Match the discussion in Section 1, Remark 1, with the $p$-notation, and recall the two cases, with $\mathbb{E}[N\ln N]<\infty$ or $\mathbb{E}[N\ln N]=\infty$. In both cases, we do have $\mathbb{E}[N_{k}]=c^{k}$.
  • For the case $\mathbb{E}[N\ln N]<\infty$, or equivalently, $\mathbb{E}[Z]=1$, we can let $p=c^{-k}$ and $\nu_{p}=N_{k}$. Then
    $\mathbb{E}[\nu_{p}]=\frac{1}{p},\quad p\nu_{p}\stackrel{{\scriptstyle d}}{{\to}}Z,\quad p\downarrow 0.$
  • If $\mathbb{E}[N\ln N]=\infty$, or equivalently, $\mathbb{E}[Z]=\infty$, then either let $p=c^{-k}$; then $\mathbb{E}[\nu_{p}]=1/p$ but not $p\nu_{p}\to Z$; or let $p=1/C_{k}$, then $p\nu_{p}\to Z$ but $\mathbb{E}[\nu_{p}]\neq 1/p$. We prefer parameterization $p=1/C_{k}$.
In the second case, we cannot have both expectation and convergence assumptions for $\nu_{p}$. The notation of $\nu_{p}$ is used in the following literature (the list is not exhaustive): [42, Section 2], (1) and (3), both expectation and convergence; [30, Section 1], only convergence; [24, page 304], they mentioned only expectation but clearly meant convergence as well, as seen from Theorem 2; [25, Section 2], only expectation, but later convergence is introduced; [30, Section 2] (3) and (4), only convergence; [13, Chapter 4], (6.32), only convergence. Sometimes, commutativity is required or derived in special cases.
The real reason behind the prevalence of this $p$-notation is that, historically, random stability study to a large extend (although not exclusively) was focused on geometric stability (that is, when $N=N_{p}$ is geometric from (11)). In this case, $pN_{p}$ converges to the standard exponential distribution as $p\downarrow 0$, and $\mathbb{E}[N_{p}]=1/p$. Approximations of sums of $N_{p}$ many independent identically distributed random variables as $p\downarrow 0$ (similar in spirit to the Central Limit Theorem) were extensively studied. A good exposition is in [23, Chapter 9].
Lemma 10.
Pick a distribution $Q$ on the half-line. Assume $G(Q)$ is nontrivial. Then the random variable $N_{c}$ with probability generating function (19) with $L$ being the Laplace transform of $Y$ satisfies
(22) $\frac{N_{c}}{c}\stackrel{{\scriptstyle d}}{{\to}}Y,\quad c\to\infty,\quad c\in S(Q),$
if and only if $\mathbb{E}[N_{c}\ln N_{c}]<\infty$. In this case, $\mathbb{E}[Y]=1$.
3.5. Explicit-form semigroups
The geometric distribution has the semigroup property: The probability generating function $G_{p}$ of $N_{p}$ satisfies
(23) $G_{p}\circ G_{q}=G_{q}\circ G_{p}=G_{pq}.$
Until recently, this geometric family (and its extensions, see below) were the only explicit families of probability generating functions $(G_{p},\,p\in(0,1])$, having finite mean $G^{\prime}_{p}(1)<\infty$, and this property (23). Extensions include the modifications of geometric distributions with atoms at zero. Their probability generating functions are general-form fractional-linear. Another extension includes the infinitely divisible case above. Then $X$ is Gamma with shape $1/k$ instead of an exponential.
Another example is the Sibuya random variable $N_{p}$ with probability generating function $G_{p}(s)=1-(1-s)^{p}$, introduced in [50] with $\beta=\gamma=1$ in (2); see also a more recent article [31] with an extension with mass at zero. Such family of distributions satisfy (23) as well. But they have infinite mean. As follows from Theorem 1, do not have any $N_{p}$-stable distributions, for any $p\in(0,1)$.
Until recently, these were the only known exact-form commuting semigroups. However, recently there were breakthroughs on this topic, with several more explicit commuting families of probability generating functions, and with finite mean greater than $1$, including a generalization of the Sibuya distribution. This corresponds to continuous-time branching processes $N=(N(t),\,t\geq 0)$ (See Section 6 for more details), with
(24) $G_{p}(s)=\mathbb{E}\left[s^{N(t)}\right],\quad p=e^{-t}.$
From (23), the composition of probability generating functions of $N(t)$ and $N(t^{\prime})$ is the probability generating function of $N(t+t^{\prime})$. Comparing (21),  (22), (24) we have the following reparametrization rule:
(25) $p=\frac{1}{c}=e^{-t},\quad c=e^{t}=\frac{1}{p}.$
4. Equivalent Definitions of Random Stability
There is another definition of random stable variables in the literature, one that is based on weak limits of scaled random sums. Just as stable and strictly stable distributions (in the classic sense) are often defined as weka limits of scaled sums of $N$ independent identically distributed random variables as $N\to\infty$, so can be done for random stable distributions. See [30, Section 2], and [38, Theorems 1, 2, 3].
Instead of $N\to\infty$, we need a semigroup of probability distributions on nonnegative integers; or, equivalently, a semigroup of probability generating functions. We might as well assume this semigroup is commutative. Although we think one does not really need this assumption, the proof is easy if we impose it.
Luckily, even when we start from one $N$ such that $X$ is $N$-stable, such semigroup comes naturally: This is the discrete semigroup generated by $N$; or, equivalently, the set of probability generating functions of $N_{k}$ for each $k$, where $(N_{k})$ is the discrete-time branching process with $N$ offsprings.
The result below is simple but needs to be included separately. We could not find it explicitly written in the literature.
Lemma 11.
Take a random variable $Y\geq 0$ with $\mathbb{E}[Y]=1$. Assume its semigroup $G(Y)$ is not trivial. Let $N_{c}$ be the random variable with probability generating function as in (19), with $c\in S(Y)$. Then a random variable $X$ is $N$-stable for at least one, and therefore for all, $N$ with distribution in $G(Y)$, if and only if there exists a sequence of independent identically distributed random variables $U_{1},U_{2},\ldots$ independent of all these $N$, and a function $a:S(Y)\to[0,\infty)$ such that
(26) $\frac{U_{1}+\ldots+U_{N_{c}}}{a(c)}\stackrel{{\scriptstyle d}}{{\to}}X,\quad c\to\infty,\quad c\in S(Y).$
If this is true, then for any such sequence $U_{1},U_{2},\ldots$ the statement (26) is equivalent to (27):
(27) $\frac{U_{1}+\ldots+U_{[c]}}{a(c)}\stackrel{{\scriptstyle d}}{{\to}}Z,\quad c\to\infty,\quad c\in S(Y).$
In this case, we have (7), with $Z$ strictly $\alpha$-stable, and $Y$ and $Z$ independent.
We see that $U_{j}$ belong to the strict domain of attraction: Recall our discussion above about defining strictly stable distributions as weak limits of scaled sums of independent identically distributed random variables. We stress this is different from classic $\alpha$-stable domains of attraction, where adding constants is allowed: Recall (6) and the discussion there. Strict domain of stability include scaling sums of independent indentically distributed random variables, but not adding constants. See also discussion in [30, Section 2].
5. Semigroup Classification
Now we start the discussion of the structure of the semigroup $S(X)\subseteq[1,\infty)$ for a given nonnegative random variable $X$ with $\mathbb{E}[X]=1$. It is time to state the main open question: Consider a distribution $Q$ on $0,\infty)$ with mean $1$. Find all distributions $P$ on $\{0,1,2,\ldots\}$ such that $Q$ is $P$-stable. Find the semigroup structure of $G(Q)$. Is it discrete or continuous? If it is discrete, is it cyclic, or not finitely generated? Can we generalize some construction techniques for the general case? Below, we provide a partial classification of such semigroups. The open problem is to complete this classification.
5.1. Trivial semigroup
$S(X)=\{1\}$ is the trivial semigroup. Examples: Lemma 7 above, or $X$ which is not absolutely continuous.
5.2. Full semigroup
$S(X)=[1,\infty)$ is the full semigroup. Example: geometric $N_{p}$ from (11), with $p\leftrightarrow 1/p$ as the isomorphism. More general discussion on this case, which is equivalent to continuous-time branching processes, is in Section 3.
5.3. Cyclic semigroup
$S(X)=\{1,a,a^{2},\ldots\}$ for some $a>1$. A construction was given in [7, Examples 2, 4] based on [9, Section 5] and [4] but in our view the proofs are not witten well, and they should be improved. Particular cases of $Q$ discussed in [4, Section 6] are as follows: The standard $P$-stable $Q$ for the following $P$: (a) $P(\{1\})=P(\{2\})=0.5$, with $\varphi(s)=(s+s^{2})/2$; (b) $P$ is the shifted geometric, with probability generating function $\varphi(s)=s^{2}/(4-3s)$, used in [2, Section 3] as a tool to construct Brownian motion on the Sierpinski gasket. The idea is to create an almost constant solution $\theta:(0,\infty)\to(0,\infty)$ to the functional equation $\theta(ax)\equiv b\theta(x)$, for fixed $a,b>0$. See [55] for additional information, also [46, Section 4].
Also, this semigroup corresponds to a discrete-time supercritical branching process $(N_{k})$ with $N$ offsprings, with $\mathbb{E}[N\ln N]<\infty$ and $\mathbb{E}[N]=a$. We consider $X=\lim(N_{k}/a^{k})$. Beyond the general statement that $X$ must be absolutely continuous on $(0,\infty)$, there was research on the behavior of $X$ at zero and infinity. There are two very distinct cases: Schröder, with $\mathbb{P}(N=0)+\mathbb{P}(N=1)>0$, and Böttcher, with $N\geq 2$ almost surely. The former case was discussed in a very recent article [33]. Both cases were studied in an early article [5]: Asymptotics at zero for Schröder case in [5, Section 2]; Asymptotics at zero for Böttcher case in [5, Section 3]; Right tail was studied in [5, Section 4]. Asymptotics at zero for the Böttcher case was also studied in [12].
A related question is embeddability: Whether the discrete-time branching process $(N_{k})$ can be represented as a skeleton for a continuous-time branching process $N(k\triangle t)$ for some $\triangle t>0$. This was studied in detail in [21], see also [1, Chapter III], Section 12. This is equivalent to upgrading from $S(Q)\supseteq\{1,a,a^{2},\ldots\}$ to $S(Q)=[1,\infty)$. This is distinct from taking a distribution $Q$ and asking whether $S(Q)=[1,\infty)$. To answer the latter question, we refer to Section 6 below.
5.4. Natural numbers
$S(X)=\{1,2,3,\ldots\}$. This is true for $X=1$, and $Q=\delta_{1}$, since $L(u)=e^{-u}$, and $L^{\leftarrow}(s)=-\ln s$, and for $c\geq 1$, we have: $L(cL^{\leftarrow}(s))=e^{c\ln s}=s^{c}$, which is a probability generating function if and only if $c=1,2,3,\ldots$, see [7, Example 3].
5.5. Complete squares
$S(X)=\{1,4,9,16,\ldots\}$. This was discussed in [24, Section 4] for $L(u)=1/\mathrm{ch}(\sqrt{2u})$, where $\mathrm{ch}(x)=(e^{x}+e^{-x})/2$ is the hyperbolic cosine, with the inverse function $\mathrm{arcch}(s)=\ln(s+\sqrt{s^{2}-1})$.
Lemma 12.
The function $\varphi_{c}(s)=L(cL^{\leftarrow}(s))$ from (19) is a probability generating function if and only if $c=n^{2}$ for positive integer $n$.
We can view $X$ as the first hitting time of $\{\pm 1\}$ by the standard Brownian motion $W=(W(t),\,t\geq 0)$ starting from zero: $X\stackrel{{\scriptstyle d}}{{=}}\inf\{t\geq 0\mid|W(t)|=1\}$. We refer to [20, Chapter 2, Section 2.8.C, (8.29)] for $x=a/2$ and $a=2n$. For $c=n^{2}$, the probability generating function is
(28) $\varphi(s)=L(n^{2}L^{\leftarrow}(s))=\frac{1}{T_{n}(1/s)},$
where $T_{n}(u)=\cos(n\arccos(u))=\mathrm{ch}(\mathrm{arcch}(u))$ is the $n$th degree Chebyshev polynomial. The probability generating function $\varphi$ in (28) corresponds to the distribution of the hitting time $\tau_{n}$ of $\{\pm n\}$ by the simple symmetric random walk $(S_{k})$, starting from 0:
$S_{k}=Z_{1}+\ldots+Z_{k},\quad\mathbb{P}(Z_{k}=\pm 1)=\frac{1}{2},\quad Z_{k}\quad\mbox{i.i.d.}$
Then $\tau_{n}=\min\{k\geq 0\mid|S_{k}|=1\}$. See [11, Chapter XIV], Section 4 for general discussion of this hitting time as the duration of gambling, and Section 5 for explicit form of the generating function with $p=q=1/2$, see (5.2) and (5.4) on page 352. We can interpret
$W_{\tau_{n}}:=\inf\{t\geq 0\mid|W(t)|=n\}\stackrel{{\scriptstyle d}}{{=}}n^{2}X.$
This equality in law follows from the self-similarity of the Brownian motion, see [20, Chapter 2]. We also refer the reader to [24, Section 3] with an interesting discussion about rational probability generating functions.
6. Continuous-Time Branching Processes
6.1. Construction
Pick a probability measure $Q$ on $0,\infty)$ with Laplace transform $L$. This is the case when $S(Q)=[1,\infty)$, or, equivalently, when for any $c\geq 1$ the function (19) is a probability generating function of some distribution on $\{0,1,2,\ldots\}$. In other words, we have a family of commuting semigroups (23) for $p\in(0,1]$; and this corresponds to a continuous-time branching process $N=(N(t),\,t\geq 0)$ via (24).
This branching process is described as follows, as outlined in [17, Chapter 5], Theorem 9.1, and [1, Chapter III], Sections 2 and 3, see also a short discussion in [51, Chapter V], Section 8. Start, as in discrete time, from $N(0)=1$. Fix a distribution $H$ on $\{0,2,3,\ldots\}$ called the generating distribution. This $H$ plays the role similar to the infinitesimal generator for diffusion processes. Each particle which exists at time $t$ is replaced during time interval $t,t+\triangle t]$ with small $\triangle t$, with probability $\triangle t$ with the number of particles distributed as $H$.
This continuous-time branching process contains skeletons: Discrete-time branching processes,
(29) $N_{k}=N(k\triangle t),\,k=0,1,2,\ldots$
Their offspring distribution has probability generating function $G_{p}$ for $p=\exp(-\triangle t)$. The converse question is embeddability: Whether a discrete-time branching process $(N_{k})$ is really a skeleton of some continuous-time branching process $(N(t),\,t\geq 0)$, as in (29), see [21] and [1, Chapter III], Section 12.
A necessary and sufficient condition for existence and non-explosion (lack of $t>0$ such that $X(t)=\infty$) is, see [17, Chapter 5, (9.3)]: The probability generating function $h$ of $\mathcal{H}$ needs to satisfy
(30) $\int_{0.5}^{1}\frac{\mathrm{d}u}{u-h(u)}=\infty.$
A particular case is $\mathbb{E}N(t)<\infty$ (equivalently, finite mean $c=h^{\prime}(1)$ of $H$, which implies (30)). In this case, scaling asymptotics holds in continuous time, [17, Chapter 5, Section 11.3]:
(31) $e^{-(c-1)t}N(t)\stackrel{{\scriptstyle d}}{{\to}}Y,\quad t\to\infty.$
For such $Y$, for any $c>1$ $L(cL^{\leftarrow}(\cdot))$ is a probability generating function; so the above semigroup is $1,\infty)$. Using the alternative notation from (25) we have
(32) $a(p)N_{p}\stackrel{{\scriptstyle d}}{{\to}}Y,\quad p\downarrow 0,\quad a(p):=p^{c-1}.$
As before, $Y\geq 0$, and we can express the inverse Laplace transform $L^{\leftarrow}$ of $Y$ as a simple function of $h$:
(33) $L^{\leftarrow}(s)=\exp\left[-\int_{s}^{1}\frac{h^{\prime}(1)-1}{h(x)-x}\,\mathrm{d}x\right]$
This gives us a way to construct many continuous commuting semigroups $(G_{p},0<p\leq 1)$ as in (23) which have the common random stable $X$. See [17, Chapter 5], Section 11, (11.7) or [1, Chapter III], Section 8, Theorem 3.
6.2. Yule and Neveu processes
For geometric distributions, $(N(t),\,t\geq 0)$ is called the Yule process, introduced by Yule in [54, Section 1] in 1925 for modeling evolution of the number of species as the pure birth process with birth rate $n$ at state $n$. See the discussion in [17, Chapter 5, Section 8] or [1, Chapter III], Section 5. For Sibuya distributions, this is called the Neveu process, introduced in an unpublished manuscript [44] by Neveu in relation to Generalized Random Energy Model (GREM) of Derrida. Links to Sherrington-Kirkpatrick model and GREM are discussed in [3, Chapter 6]. The exponential $\mathrm{Exp}(1)$ from (13) is $G_{p}$-stable for any $p$. However, there is no Sibyua-stable random variable, since the mean of the Sibyua distribution is zero. Properties of the Neveu process are discussed in [18]. They also considered generalizations with Sibuya distribution with an atom at zero in [18]. These branching processes have nonzero extinction probability.
For the Yule process with geometric distributions, the generating distribution is
$\mathbb{P}(H=2)=1,\quad h(s)=s^{2}.$
For the Neveu process with Sibuya distribution, the generating distribution is
$\mathbb{P}(H=n)=\frac{1}{n(n+1)},\quad n=2,3,\ldots,\quad h(s)=s+(1-s)\ln(1-s),$
see also [18, Subsection 1.3, formula (16)]. The result (31) applied to the Yule process with geometric distributions states
$e^{-t}N(t)\stackrel{{\scriptstyle d}}{{\to}}Y\sim\mathrm{Exp}(1),\quad t\to\infty;$
Using the equivalent notation as in (25), the result (32) applied to the Yule process states:
$p\Gamma_{p}\stackrel{{\scriptstyle d}}{{\to}}Y\sim\mathrm{Exp}(1),\quad p\downarrow 0.$
Here, $c=2$, since the mean of the constant random variable 2 is clearly 2. For the Neveu process with infinite mean, a very different asymptotics holds:
$e^{-t}\ln N(t)\to\mathrm{Exp}(1),\quad t\to\infty;$
or, equivalently, after reparametrization as in (25):
$p\ln G_{p}\to\mathrm{Exp}(1),\quad p\downarrow 0,$
See the original proof in [8] for the discrete time (in their notation, $F(s)=1-(1-s)^{e^{-t}}$ and $g(s)=s^{e^{t}}$ so $a=1$ and $b=e^{t}-1$) and [14, Theorem 3] for the continuous time ($\alpha=1,q=0$ in their notation).
There exist many other then Yule continuous-time branching processes. Equivalently, there exist many other than geometric families satisfying (16). This contradicts the previous statement [29, Remark 2.4] by T. Kozubowski. All these new continuous-time branching processes below are intermediate between the Yule process (one particle can only create two at a time) and the Neveu process (one particle can create arbitrarily many, and this random number has infinite mean).
6.3. Generating geometric distribution
Here we present a special case of geometric generating $H$. The full treatment is given in a recent article [40], with complicated formulas for mass functions and PGFs, using special functions. Our example is easily treatable and explicit. We also refer to another article [41] by the same authors, treating the case of Poisson $H$. This involves complicated special functions, too. Similar computations were done for [39, Theorem 1] by the same authors; but they consider the subcritical case $c<1$, with a.s. extinction; of course, this branching process does not have a limiting $X$. One possible application is to physics: the Bose-Einstein distribution have geometric as one possible limit, see [19, (9)], and this can be connected to the current process.
Take the generating distribution $\mathbb{P}(H=n)=2^{1-n}$ for $n=2,3,\ldots$ Shifted geometric with parameter $p=0.5$, with PGF $h(s)=0.5s^{2}(1-0.5s)^{-1}$. This $H$ generates a continuous-time branching process with finite mean $\mathbb{E}[N(t)]=e^{2t}$ at time $t$. Its PGF can be found in exact form:
$\psi_{t}(s)=\frac{2s}{s^{2}+4(1-s)e^{2t}},\quad t\geq 0,\,s\in[0,1].$
The limiting distribution $Y$ from (33) is shifted Mittag-Leffler (or Kovalenko) with index $1/2$ and the following Laplace transform:
$\mathbb{E}[e^{-uY}]=\frac{2}{1+\sqrt{1+4u}},\quad u\geq 0;$
There is yet another case of quite explicit formulas for composition semigroups, in the case of geometric $H$ modified at zero; and Poisson $H$. These are related to Lambert $W$-function, the solution to the equation $W(x)e^{W(x)}=x$. We will not present details here, instead referring readers to recent literature [39, 40, 41]. They have finite mean and nonzero extinction probability.
6.4. Theta-branching processes
In a recent article [34], they introduced a classification of 9 cases, see Section 4, with generalization of a Neveu process with Sibuya distribution. It has four parameters, including $\theta\in[-1,1]$. The case 4 is precisely the Sibuya distribution modified so that there is an atom at zero, and a nonzero extinction probability. For us, the case 3 is the most interesting, since it corresponds to the supercritical case, but with finite mean. The probability generating function $h$ of the generating distribution $H$ is, see [34, Section 7]:
$h(s)=s+\frac{(1-s)^{1+\theta}-(1-q)^{\theta}(1-s)}{1+\theta-(1-q)^{\theta})}.$
The probability generating function of the continuous-time branching process at time $t$ with $p=e^{-t}$, see [34, Section 4]:
$G_{p}(s)=1-\left[p(1-s)^{-\theta}+(1-p)(1-q)^{-\theta}\right]^{-1/\theta}.$
See [15, Proposition 7, page 1087] for the distribution of the standard $N$-stable random variable $Y$ from (33).
7. Proofs
In this section, first we present proofs of three main results: Theorems 123. Next, we present proofs of the twelve lemmas. Finally, we state and prove a technical lemma from complex analysis.
7.1. Proof of Theorem 1
This is the longest proof, and we present a short overview. We split it in 8 steps. Step 1 is introductory and involves branching processes and random sums of copies of $X$. Step 2 is $\mathbb{E}[N]\leq 1$. Steps 3–7 are for the case $\mathbb{E}[N]=\infty$. We symmetrize the distribution $X$ and then use tightness arguments to arrive at a contradiction. Finally, Step 8 mentions the classic case $1<\mathbb{E}[N]<\infty$.
Step 1. As before, let $\varphi$ be the probability generating function of $N$. Consider a discrete-time branching process $(N_{k})$ starting from $N_{0}=1$ and each particle having the number of offsprings distributed as $N$. The probability generating function of $N_{k}$ is the $k$th composition of $\varphi$:
$\varphi_{k}(s)=\mathbb{E}\left[s^{N_{k}}\right]=\varphi(\varphi(\ldots\varphi(s)\ldots)).$
For background, we refer to the classic monographs [17, Chapter 1] and [1, Chapter I]. Applying the main equality $k$ times, we get:
(34) $X_{1}+\ldots+X_{N_{k}}\stackrel{{\scriptstyle d}}{{=}}c^{k}Z_{1}.$
This identity is helpful for the proofs of lack of $N$-stable $X$ for $\mathbb{E}[N]\geq 1$ and $\mathbb{E}[N]=\infty$. The first case is simple, the second is much harder. We always assume $(N_{k})$ is independent of $X_{1},X_{2},\ldots$
Step 2. Assume $\mathbb{E}[N]\leq 1$. From the classic theory of branching processes [17, Chapter 1], Theorem 6.1 or [1, Chapter I], Section 5, Theorem 1, the process $(N_{k})$ becomes extinct with probability $1$. THat is. $N_{k}=0$ for some $k$ almost surely. This immediately implies that the left-hand side of (34) is equal to zero. Therefore, the right-hand side of (34) is also zero, which implies $X_{1}=0$ almost surely. This contradiction completes the proof of lack of nontrivial $N$-stable $X$ if $\mathbb{E}[N]\leq 1$.
Step 3. Now assume $\mathbb{E}[N]=\infty$. Define $Y_{n}:=X_{n}-X^{\prime}_{n}$, where $X^{\prime}_{1},X^{\prime}_{2},\ldots$ is yet another sequence of copies of $X$, independent of each other, of $X_{1},X_{2},\ldots$ and of the branching process $(N_{k})$. Then
(35) $Y_{1}+\ldots+Y_{N_{k}}=\left(X_{1}+\ldots+X_{N_{k}}\right)-\left(X^{\prime}_{1}+\ldots+X^{\prime}_{N_{k}}\right).$
Dividing (35) by $c^{k}$, we get:
(36) $\Sigma_{k}$ $:=\frac{1}{c^{k}}\left(Y_{1}+\ldots+Y_{N_{k}}\right)=S_{k}-S^{\prime}_{k},$
(37) $S_{k}$ $:=\frac{1}{c^{k}}\left(X_{1}+\ldots+X_{N_{k}}\right),\quad S^{\prime}_{k}=\frac{1}{c^{k}}\left(X^{\prime}_{1}+\ldots+X^{\prime}_{N_{k}}\right).$
Note that $S_{k}$ and $S^{\prime}_{k}$ are dependent via $N_{k}$. Thus we cannot claim that $\Sigma_{k}\stackrel{{\scriptstyle d}}{{=}}X-X^{\prime}$ for an independent copy $X^{\prime}$ of $X$. However, we can claim the sequence $(\Sigma_{k})$ is tight (in other words, relatively compact, bounded in probability): For any $\varepsilon>0$, there exists a $K>0$ large enough so that $\mathbb{P}(|\Sigma_{k}|\geq K)\leq\varepsilon$ for all $k$. This follows from the tightness of sequences $(S_{k})$ and $(S^{\prime}_{k})$, which all have the same distribution, the same as $X$.
Step 4. Define the characteristic function of $\Sigma_{k}$:
(38) $g_{k}(u)=\mathbb{E}\left[e^{\mathrm{i}u\Sigma_{k}}\right]$
It is real-valued, since $\Sigma_{k}\stackrel{{\scriptstyle d}}{{=}}-\Sigma_{k}$. It turns out that this simplifies the proof in a critical way. See [36, Chapter 3], Theorem 3.1.2. Applying [52, Chapter 3], Lemma 3.1.3, we derive from tightness of the sequence $(\Sigma_{k})$ that
(39) $\sup\limits_{k\geq 1}|1-g_{k}(u)|\to 0,\quad u\to 0.$
Since $g_{k}$ is real-valued, and by [36, Chapter 2], page 36, $g_{k}(u)\leq 1$, we can simply write this statement (39) without the absolute value. This characteristic function $g_{k}$ from (38) can be represented as
(40) $g_{k}(u)=\varphi_{k}(g(u/c^{k})),\quad g(u):=\mathbb{E}\left[e^{\mathrm{i}uY_{n}}\right]=\mathbb{E}\left[e^{\mathrm{i}u(X_{n}-X^{\prime}_{n})}\right]=|f(u)|^{2}$
is the characteristic function of $Y_{n}$, see [36, Chapter 3], Corollary 2 of Theorem 3.3.1.
Step 5. It follows from scaling theory for discrete-time branching processes, see for example [48, Theorem 4.4] with infinite mean that for a fixed $a>1$, the event $A=\{N_{k}/a^{k}\to\infty\}$ has positive probability. For any $s\in(0,1)$, we can rewrite
$A=\left\{s^{N_{k}/a^{k}}\to 0,\quad k\to\infty\right\}$
and note $0\leq s^{N_{k}/a^{k}}\leq 1$ almost surely. Applying the Fatou lemma, we get:
$\varphi_{k}(s/a^{k})=\mathbb{E}\left[s^{N_{k}/a^{k}}\right]$
satisfies the upper limit
[…]
Step 6. Comparing (41) with (39), we get: For any $u,s\in(0,1)$ and $a>1$, there exists a $k_{0}(u,s,a)$ such that for $k\geq k_{0}(u,s,a)$ we have:
(42) $g(u/c^{k})\geq s/a^{k}.$
Take the logarithms in (42) and apply the elementary inequality $\ln y\leq y-1$ for $y>0$:
(43) $g(u/c^{k})-1\geq a^{k}\ln s.$
Dividing (43) by $a^{k}$ and letting $k\to\infty$, we get:
[…]
In the right-hand side of (44), the number $s\in(0,1)$ is arbitrary. Therefore, we can take $s$ as close to $1$ as we wish, to make $\ln s$ negative but as close to zero as we wish. Also, recall that by [36, Chapter 2], page 36, we have: $g(uc^{-k})\leq 1$. Applying all this to (44), we get:
(45) $\lim\limits_{k\to\infty}a^{-k}(g(uc^{-k})-1)=0.$
Step 7. Rewrite as second-order discrete difference with step $t_{k}:=0.5c^{-k}u$:
$2(g(uc^{-k})-1)=g(-uc^{-k})+g(uc^{-k})-2g(0)=\triangle^{t_{k}}_{2}g(0)$
where we define $\triangle^{t}_{2}g(u):=g(u-2t)+g(u+2t)-2g(u)$. Clearly, $t_{k}\to 0$, and letting $a=c^{2}$ (since $a$ is arbitrary), we rewrite (45) as $\triangle_{2}^{t_{k}}g(0)/t_{k}^{2}\to 0$. Apply [36, Chapter 2], Theorem 2.3.1 and conclude: $g^{\prime\prime}(0)=0$, thus $\mathbb{E}[Y^{2}]=0$, and $Y=0$ almost surely, thus $g(u)\equiv 1$. Comparing this with (40), we get: $|f(u)|\equiv 1$. Apply [36, Chapter 2], Theorem 2.1.4, Corollary 1. We get: $f(u)=e^{\mathrm{i}uX_{0}}$ and $X=x_{0}$ almost surely for some $x_{0}\in\mathbb{R}$. But $X$ is $N$-stable. Plugging $X=x_{0}$ into (1), we get: $x_{0}N=cx$. Thus either $N=c$, which contradicts $\mathbb{E}[N]=\infty$; or $x_{0}=0$, which implies $X=0$ almost surely, but we exclude this trivial case. This completes the proof that there is no $N$-stable $X$ in case $\mathbb{E}[N]=\infty$.
Step 8. The classic case $1<\mathbb{E}[N]<\infty$ is well-known and discussed in the Introduction, using scaling limits of dsicrete-time supercritical branching processes. See [1, Chapter I], Section 10, Theorems 2 and 3.
7.2. Proof of Theorem 2
Step 1. We show that any random variable $X$ with characteristic function $f=L\circ g$ and $g\in\mathcal{G}$ is $N$-stable. Apply (9) and (2) with $c=\mathbb{E}[N]$:
$\varphi(f(u))=\varphi(L(g(u)))=L(cg(u))=L(g(c^{1/\alpha}u))=f(c^{1/\alpha}u).$
Incidentally, this proves that characteristic functions of the left- and right-hand sides of (1) coincide. Thus we have equality in law.
Step 2. We show that any random variable $X$ with characteristic function
$f(u)=\mathbb{E}\left[e^{\mathrm{i}uX}\right]=L(g(u))$
for some $g\in\mathcal{G}$ of index $\alpha$ can be represented as $X=Y^{1/\alpha}Z$, where $Y$ is the standard $N$-stable random variable, and $\mathbb{E}\left[e^{\mathrm{i}uZ}\right]=e^{-g(u)}$. We simply apply [30, Proposition 3.1] and use their remarks on page 309, (10) – (12), Remark 1, about the difference between stable and strictly stable. One can also consult [45, Chapter 1], Section 2 for the latter question.
Step 3. Take a characteristic function $f$ of an $N$-stable random variable $X$. We need to prove that this solution to the main equation (2) can be represented as $f=L\circ g$ for some $g\in\mathcal{G}$. We apply results [13, Chapter 4], Section 4.6. Recall the discussion about parameterization. Match the notation:
$\Theta=\{c^{-1},c^{-2},\ldots\},\quad\nu_{c^{-k}}=N_{k},\quad\nu=X.$
As in the proof of Theorem 1, we conclude that if $X$ is $N$-stable, then $X_{1}+\ldots+X_{N_{k}}\stackrel{{\scriptstyle d}}{{=}}c^{k}X_{1}$. Rewrite this using our new notation:
$X_{\theta,1}+\ldots+X_{\theta,N_{k}}\stackrel{{\scriptstyle d}}{{=}}X,\quad\theta=c^{-k},\quad X_{\theta,j}:=\frac{X_{j}}{c^{k}}.$
In the notation of [13, Chapter 4], we have (6.33) with $F$ being the CDF of $X$. Actually, our statement is even stronger: We have equality in law, not just convergence in law. Also, we have (6.32) in the same notation. This was discussed in the subsection on parameterization. Therefore, by [13, Chapter 4], Theorem 4.6.3, the formula (6.34) holds with an infinitely divisible $Y$ with the CDF $G$ and with $m(\theta)=[1/\theta]=[c^{k}]$. Rewrite this conclusion in our original notation:
$\frac{X_{1}+\ldots+X_{[c^{k}]}}{c^{k}}\stackrel{{\scriptstyle d}}{{\to}}Y.$
By the classic theory of stable distributions, this implies $Y$ is strictly stable. And thus $\mathbb{E}\left[e^{\mathrm{i}uY}\right]=e^{-g(u)}$ for some $g\in\mathcal{G}$. By [13, Chapter 4], Theorem 4.6.5, we can write (6.30), which completes the proof of the representation $f=L\circ g$. With this, we proved that a random variable $X$ is $N$-stable if and only if its characteristic function $f$ can be represented as $L\circ g$ with $g\in\mathcal{G}$, and thus completed the proof of Theorem 2.
7.3. Proof of Theorem 3
Step 1. From Theorem 2, the characteristic function $f$ of $X$ can be represented as $f(u)=L(g(u))$, where $L$ is the Laplace transform of the standard $N$-stable $Y$, and $g\in\mathcal{G}$. We also have (2) with $c=\mathbb{E}[N]$. Finally, $\mathbb{E}[Y]=1$.
Step 2. Assume $M$ and $N$ commute. Then their probability generating functions $\varphi$ and $\psi$ satisfy (16), and therefore
(46) $\varphi(\psi(L(u)))=\psi(\varphi(L(u)))=\psi(L(cu)).$
The function $\psi\circ L$ is the Laplace transform of the random variable $S=Y_{1}+\ldots+Y_{M}$, where $Y_{1},Y_{2},\ldots$ are copies of $Y$ independent of each other and of $M$. From (46), we see that $S$ is $N$-stable, and has finite mean $\mathbb{E}[S]=\mathbb{E}[Y]\cdot\mathbb{E}[M]=\mathbb{E}[M]=b$. By the uniqueness of the standard $N$-stable random variable, $S/b\stackrel{{\scriptstyle d}}{{=}}Y$. If we write this equality in law in terms of Laplace transforms, we have: $\psi(L(u))=L(bu)$. Therefore, $Y$ is $M$-stable. Finally, letting $\alpha$ be the index of $g$, we get:
$\psi(f(u))=\psi(L(g(u)))=L(bg(u))=L(g(b^{1/\alpha}u))=f(b^{1/\alpha}u).$
This proves $X$ is $M$-stable as well.
Step 3. Conversely, if $X$ is both $M$-stable and $N$-stable, then $\varphi(f(u))=f(au)$ and $\psi(f(u))=f(bu)$. Applying both these identities, we get:
$(\psi\circ\varphi)(f(u))=f(abu)=(\varphi\circ\psi)(f(u)).$
If $X$ is a constant, then $M$ and $N$ are also constants, and obviously they commute. If $X$ is not a constant, then $f$ takes at least one value $z$ inside the unit disc $\mathbb{D}=\{z\in\mathbb{C}\mid|z|<1\}$, [36, Chapter 2], Theorem 2.1.4, Corollary 2. But $f$ is continuous, therefore this value $z$ is the limit point of the image of $f$. Both $\varphi\circ\psi$ and $\psi\circ\varphi$ are probability generating functions of some distributions. Therefore, they are analytic on $\mathbb{D}$. Applying the classic result from compelx analysis, we get (16).
Step 4. If $X$ is $N$-stable and $M$-stable, then we have: $\mathbb{E}[M]\in(1,\infty)$ and $\mathbb{E}[M\ln M]<\infty$; also $\mathbb{E}[N]\in(1,\infty)$ and $\mathbb{E}[N\ln N]<\infty$.
Step 5. Finally, the last claim: For distinct $M$ and $N$, the representation $f=L\circ g$ has the same $L$ and the same $g$. This follows from the uniqueness Lemma 1. In particular, the index $\alpha$ for both $M$ and $N$ is the same.
7.4. Proof of Lemma 1
Step 1. It is clear from the representation of $g_{j}\in\mathcal{G}$ in (4):
(47) $g_{j}(u)=|u|^{\alpha_{j}}(\beta_{j}+\gamma_{j}\mathrm{i}\mathrm{sgn}(u))$
that $\mathrm{Re}g_{j}(u)>0$ for $u\neq 0$; and $g_{j}(u)\to 0$ as $u\to 0$. And by Lemma 54
(48) $\frac{1-L_{j}(g_{j}(u))}{g_{j}(u)}\to 1,\quad j=1,2,\,u\to 0.$
Indeed, $L_{j}$ is the Laplace transform of a probability measure on $0,\infty)$ with mean $1$. Next,
(49) $L_{1}(g_{1}(u))=L_{2}(g_{2}(u)).$
Comparing (48) with (49), we get: $g_{1}(u)/g_{2}(u)\to 1$ as $u\to 0$.
Step 2. From (47), rewrite for $u>0$:
(50) $\frac{g_{1}(u)}{g_{2}(u)}=u^{\alpha_{1}-\alpha_{2}}\cdot\frac{\beta_{1}+\gamma_{1}\mathrm{i}}{\beta_{2}+\gamma_{2}\mathrm{i}}.$
Since $\beta_{1},\beta_{2}>0$, then
$D:=\left|\frac{\beta_{1}+\gamma_{1}\mathrm{i}}{\beta_{2}+\gamma_{2}\mathrm{i}}\right|>0.$
Therefore, if $\alpha_{1}>\alpha_{2}$, then apply the absolute value to (50):
(51) $\left|\frac{g_{1}(u)}{g_{2}(u)}\right|=u^{\alpha_{1}-\alpha_{2}}\cdot D\to 0,\quad u\downarrow 0.$
This statement (51) contradicts $g_{1}(u)/g_{2}(u)\to 1$. Similarly, the same happens in $\alpha_{1}<\alpha_{2}$.
Step 3. Thus $\alpha_{1}=\alpha_{2}$, and (50) becomes
$\frac{g_{1}(u)}{g_{2}(u)}=\frac{\beta_{1}+\gamma_{1}\mathrm{i}}{\beta_{2}+\gamma_{2}\mathrm{i}}.$
Letting $u\to 0$ again, we see: This number must be equal to one, therefore, $\beta_{1}=\beta_{2}$, $\gamma_{1}=\gamma_{2}$, and $g_{1}\equiv g_{2}=:g$. To complete the final step and show $L_{1}\equiv L_{2}$, recall $L_{1}(g_{1}(u))\equiv L_{2}(g_{2}(u))$. The functions $L_{1}$ and $L_{2}$ are analytic on the half-space $\mathbb{H}:=\{z\in\mathbb{C}\mid\mathrm{Re}(z)>0\}$, by Lemma 54. And the set $\{g(u)\mid u>0\}$ has limit points in $\mathbb{H}$. By the classic uniqueness result from complex analysis, $L_{1}\equiv L_{2}$. This complets the proof of Lemma 1.
7.5. Proof of Lemma 2
This is another immediate consequence of Theorem 2:
$X\stackrel{{\scriptstyle d}}{{=}}-X\Rightarrow Z\stackrel{{\scriptstyle d}}{{=}}-Z.$
The function $g$ from (4) is even: g(u)=g(u)g(u)=g(-u). This happens if and only if $\gamma=0$, so $g(u)=\beta|u|^{\alpha}$.
7.6. Proof of Lemma 3
From Theorem 2, we get: $X=\sqrt{Y}Z$, where $Z$ is strictly $2$-stable and therefore Gaussian. Thus $\mathbb{E}[X^{2}]=\mathbb{E}[Y]\cdot\mathbb{E}[Z^{2}]<\infty$, since $\mathbb{E}[Y]=1$, which follows from $\mathbb{E}[N\ln N]<\infty$.
7.7. Proof of Lemma 10
Since $X\geq 0$, by assumption, then the strictly stable random variable $Z$ with index $\alpha$ must be nonnegative as well. But this could be true only if $\alpha=1$ and $Z$ is a positive constant; or $\alpha\in(0,1)$, and $Z$ has Laplace transform (10). This can be found in [51, Chapter V], Theorem 3.5, or in the classic reference [45]. Then the Laplace transform of $X$ is
$\mathbb{E}[e^{-uX}]=\mathbb{E}\bigl[e^{-uY^{1/\alpha}Z}\bigr]=\mathbb{E}\mathbb{E}\left[e^{-uY^{1/\alpha}Z}\mid Y\right]=\mathbb{E}\left[e^{-\beta u^{\alpha}Y}\right]=L(\beta u^{\alpha}).$
7.8. Proof of Lemma 5
Using the notation from the beginning of this article, we see that the PGF of $M$ is $\varphi^{1/k}$. Therefore, the PGF of $kM$ is $\psi(s)=\varphi^{1/k}(s^{k})$. The Laplace transform of $Y$ is $L^{1/k}$. It is straightforward to check that (2) holds if and only if $\psi(L^{1/k}(u))=L^{1/k}(cu)$ for $u\geq 0$. This completes the proof.
7.9. Proof of Lemma 6
This follows from two consecutive applications of Theorem 3. Indeed, if $X$ is $N$-stable and $M$-stable, then $M$ and $N$ commute. But $Y$ is $N$-stable, and therefore $Y$ is $M$-stable.
7.10. Proof of Lemma 7
We express $L(u)=(1+u)^{-\alpha}$ for the Gamma random variable $X$ with shape $\alpha$ and scale $1$. We find the inverse function $L^{\leftarrow}(s)=s^{-1/\alpha}-1$. For $c>1$, we have from (19):
$\varphi(u)=u\left(u^{1/\alpha}(1-c)+c\right)^{-\alpha}.$
It has Taylor decomposition which contains $u^{1+1/\alpha}$. This exponent is an integer if and only if $\alpha=1/n$ for $n=1,2,3,\ldots$
7.11. Proof of Lemma 8
Step 1. If $1$ is the limit point of $S(Q)$, then for every interval $(a,b)\subseteq(1,\infty)$, however small, we have a $c\in S(P)\cap(1,b/a)$. Then for some $n$, we have $c^{n}\in(a,b)$. This $c^{n}\in S(Q)$, which proves $S(Q)$ is dense in $1,\infty)$: It is intersecting with any interval. Since $S(Q)$ is topologically closed, $S(Q)=[1,\infty)$.
Step 2. All other claims about $S(Q)$ and $G(Q)$ directly follow from [7, Proposition 1]. They use the notation $e^{t}=c$ where $t\geq 0$ and $c\geq 1$, as discussed in (25) but this is equivalent.
7.12. Proof of Lemma 9
This was, in fact, already proven in the literature, since the tails of standard $N$-stable $Y$ are related to tails of $N$: It is shown in [1, Chapter I], Section 10, Theorem 2, that for any $a\geq 0$,
$\mathbb{E}[N\ln^{1+a}(N)]<\infty\Leftrightarrow\mathbb{E}[Y\ln^{a}Y]<\infty.$
Also, it is shown in [6, Theorem 0] that for $k=2,3,\ldots$, we have:
$\mathbb{E}[N^{k}]<\infty\Leftrightarrow\mathbb{E}[Y^{k}]<\infty.$
7.13. Proof of Lemma 10
Step 1. Assume $\mathbb{E}[Y]=1$. The Laplace transform of $N_{c}/c$ is
(52) $\mathbb{E}\left[e^{-N_{c}u/c}\right]=L\left(cL^{\leftarrow}\left(e^{-u/c}\right)\right).$
If $Y$ has mean $1$, then $L^{\prime}(0+)=-1$. Derivative of the inverse function: $(L^{\leftarrow})^{\prime}(1-)=-1$. Also, an elementary calculus result shows:
$e^{-u/c}-1\sim\frac{-u}{c},\quad c\to\infty.$
Combining these asymptotics, we get:
(53) $cL^{\leftarrow}\left(e^{-u/c}\right)\to(L^{\leftarrow})^{\prime}(1-)\cdot c\cdot\left(-\frac{u}{c}\right)=u.$
Applying (53) to (52), we get that the left-hand side of (52) converges to $L(u)$, which is the Laplace transform of $Y$. This completes the if part.
Step 2. Conversely, if indeed there is such convergence, it holds for $c=b^{n}$ as well, where $b\in S(Q)$. But this corresponds to scaling of a discrete-time branching process. Using the aforementioned discussion about parameterization, we see that such scaling works only when $\mathbb{E}[Y]<\infty$, which is equivalent to $\mathbb{E}[Y]=1$.
7.14. Proof of Lemma 11
Step 1. Let us derive  (1) from (26). Direct application of [13, Chapter 4], Theorem 4.6.5, with the following notation mathc: $\theta=1/c$ and $m(\theta)=[c]$, $X_{\theta,j}=U_{j}/a(c)$, $F$ is the CDF of $X$, and $G$ is the CDF of $Z$, and $N=Y$ in (6.32), and (6.30) from [13, Theorem 4.6.3] can be rewritten as (7). See our discussion of notation above.
Step 2. Conversely, if $X$ is $N$-stable, then we can simply take $U=X$. Then we have equality in law in (26) instead of weak convergence.
Step 3. Equivalence of (26) and (27) also follows from [13, Chapter 4], Theorem 4.6.5.
7.15. Proof of Lemma 12
We compute $L^{\leftarrow}(s)=(\mathrm{arcch}(1/s))^{2}/2$. Therefore, using (19):
$\varphi(s)=L(cL^{\leftarrow}(s))=\frac{1}{\mathrm{ch}(\sqrt{c}\cdot\mathrm{arcch}(1/s))}.$
As $s\downarrow 0$, this has asymptotics $\varphi(s)=2^{1-\sqrt{c}}s^{\sqrt{c}}$. This function is analytic if and only if $\sqrt{c}=n$ is an integer, that is, $c\in\{1,4,9,\ldots\}$.
7.16. Analytical Laplace transform
Take a random variable $X\geq 0$. Consider its Laplace transform $L(u)=\mathbb{E}[e^{-uX}]$.
Lemma 13.
Assume $\mathbb{E}[X]<\infty$. We can define $L$ as a complex-valued function on the half-plane $\mathbb{H}=\{z\in\mathbb{C}\mid\mathrm{Re}(z)>0\}$. This function is analytic on $\mathbb{H}$, and we have:
(54) $\frac{1-L(z)}{z}\to\mathbb{E}[X],\quad z\to 0,\quad\mathrm{Re}(z)>0.$
Proof.
For $z=u+\mathrm{i}v\in\mathbb{H}$, we write $|e^{-zX}|=e^{-uX}$. Therefore, $\mathbb{E}|e^{-zX}|=\mathbb{E}[e^{-uX}]<\infty$, and $\mathbb{E}[e^{-zX}]$ is well defined. It is complex analytic on $\mathbb{H}$. To this end, we need to prove in the complex analytic sense: $L^{\prime}(z)=\mathbb{E}[-Xe^{-zX}]$. But this, in turn, can be proved as follows. Take two points $z,z^{\prime}\in\mathbb{H}$ and draw a segment between them. It is parametrized as $w_{t}=zt+z^{\prime}(1-t)$. We need to show
(55) $L(w_{t})-L(z)=\int_{[z,w_{t}]}\mathbb{E}\left[-Xe^{-uX}\right]\,\mathrm{d}u.$
This integral over the segment is understood in the complex analytic sense. This can be written using real-valued integration as
$\int_{0}^{t}\mathbb{E}\left[-Xe^{-(zs+(1-s)z^{\prime})X}\right](z^{\prime}-z)\,\mathrm{d}s.$
Remove expectations for a moment. From complex analysis, we get:
$e^{-(zt+z^{\prime}(1-t))X}-e^{-ztX}=\int_{0}^{t}\left(-Xe^{-w_{s}X}\right)(z^{\prime}-z)\,\mathrm{d}s.$
We need only to show that we need an interchange of integration and expectation. This requires us to use the Fubini theorem. Usually, this theorem is stated for real-vaued functions, but it works equally well for complex-valued functions. We need:
(56) $\mathbb{E}\int_{0}^{t}\left|-Xe^{-w_{s}X}\right|(z^{\prime}-z)\,\mathrm{d}s<\infty.$
Assuming $\mathrm{Re}(z)\leq\mathrm{Re}(z^{\prime})$ without loss of generality, $\mathrm{Re}(z)\leq\mathrm{Re}(w_{s})$ for all $s\in[0,1]$. Hence
$\mathbb{E}|-Xe^{-w_{s}X}|=\mathbb{E}|Xe^{-\mathrm{Re}(z)X}|<\infty.$
This proves (56), and with it proves (55). Thus $L$ is analytic on $\mathbb{H}$. The property (54) can be shown similarly, with $z=0$, since $\mathbb{E}[Xe^{-zX}]=\mathbb{E}[X]$ for $z=0$. ∎
8. Conclusions and Open Problems
This article is an attempt to review and compare various streams of existing literature on branching processes and random stability. We added a few results of our own, which we could not find in the literature; therefore this article cannot really be considered merely a review, but a hybrid between the original research and literature survey.
A natural direction for further research is: For each probability measure $Q$ on $0,\infty)$ with mean $1$, find the semigroup $G(Q)$, and describe the distributions in this semigroup.
Another research direction is subtle: Extend our results to the case $\mathbb{E}[N]<\infty$ but $\mathbb{E}[N\ln N]=\infty$. In this article, we stressed the importance of the condition $\mathbb{E}[N\ln N]<\infty$, which is not emphasized in some literature.
Yet another possible extension is smoothing transformation: For other random variables $A_{1},A_{2},\ldots$ independent of each other and of $X_{1},X_{2},\ldots$
$A_{1}X_{1}+\ldots+A_{N}X_{N}\stackrel{{\scriptstyle d}}{{=}}X.$
It reduces to random stability if $A_{j}=1/c$, and has other applications. We referred to this at the very start of this article: [35]. See also an earlier article [10].
Acknowledgements
We thank our departmental colleague Tomasz Kozubowski for raising the question whether there exists a Sibuya-stable random variable, and another question whether there exist non-geometric and non-Sibuya commuting semigroups; and further useful discussion. We thank Thierry Huillet for pointing out recent articles on continuous-time branching processes with explicit distributions. We thanks Svetlozar Rachev for multiple useful comments on an earlier draft. We thank Lev Klebanov for useful discussion, and for providing his book [22].